### v1

@@ Abstract
Abstract: Role-playing language agents (RPLAs) have emerged as promising applications of large language models (LLMs). However, simulating established characters presents a challenging task for RPLAs, due to the lack of authentic character datasets and nuanced evaluation methods using such data. In this paper, we present CoSER, a collection of a high-quality dataset, open models, and an evaluation protocol towards effective RPLAs of established characters. The CoSER dataset covers 17,966 characters from 771 renowned books. It provides authentic dialogues with real-world intricacies, as well as diverse data types such as conversation setups, character experiences and internal thoughts. Drawing from acting methodology, we introduce given-circumstance acting for training and evaluating role-playing LLMs, where LLMs sequentially portray multiple characters in book scenes. Using our dataset, we develop CoSER 8B and CoSER 70B, i.e., advanced open role-playing LLMs built on LLaMA-3.1 models. Extensive experiments demonstrate the value of the CoSER dataset for RPLA training, evaluation and retrieval. Moreover, CoSER 70B exhibits state-of-the-art performance surpassing or matching GPT-4o on our evaluation and three existing benchmarks, i.e., achieving 75.80% and 93.47% accuracy on the InCharacter and LifeChoice benchmarks respectively. Our code, dataset and models are available at https://github.com/Neph0s/CoSER.
@@ 1 Introduction
In this paper, we introduce CoSER, a collection of authentic character data, along with open state-of-the-art models and evaluation protocol based on such data, for Coordinating LLM-Based Persona Simulation of Established Roles. The CoSER dataset is sourced from narratives and dialogues in 771 renowned books, processed via our LLM-based pipeline. CoSER differs from existing datasets in two fundamental ways: 1) CoSER extracts authentic, multi-character dialogues from acclaimed literary works, in contrast to LLM-synthesized question-answer pairs in previous work. Hence, our dataset maintains high source fidelity while exhibiting greater quality and complexity. 2) CoSER incorporates comprehensive types of data, as shown in Fig.  1: i) Besides character profiles and dialogues, CoSER encompasses plot summaries, character experiences, and conversation backgrounds, supporting various purposes including prompting, retrieval, model training and evaluation. ii) Conversations in CoSER capture characters’ actions and internal thoughts beyond surface-level speech, enabling RPLAs to simulate sophisticated cognitive and behavioral processes of humans, such as “[I’m nervous, but we have to do this] (Takes a deep breath) Alright, we …”. We provide a clear comparison between CoSER and existing datasets in Table  5.
@@
    Results of extensive experiments demonstrate the significant value of our dataset for the training, retrieval and evaluation of RPLAs. Notably, the CoSER models achieve state-of-the-art performance across four benchmarks for RPLAs.
@@ 2 Related Work
An RPLA leverages LLMs to create a simulated persona $\pi_{c}$ that emulates a real character $c$ based on its persona data $\mathcal{D}_{c}$. Effective RPLAs require both comprehensive, high-quality data $\mathcal{D}_{c}$ and advanced role-playing LLMs. Among various persona types, we focus on RPLAs for established characters, which should faithfully align with their characters’ complex backgrounds and nuanced personalities.
Datasets for RPLAs  Persona data $\mathcal{D}_{c}$ describe the real persona $c$ through various representations, including profiles (Yuan et al., 2024), dialogues (Wang et al., 2024a), experiences (Li et al., 2023) and multimodal information (Dai et al., 2024), e.t.c.. As shown in Table  5, existing datasets have several limitations. 1) Many are synthesized via LLMs’ responses to general instruction sets (Wang et al., 2024a) or character-specific questions (Shao et al., 2023), such as RoleBench (Wang et al., 2024a). However, LLM-synthesized data compromise authenticity and fidelity to original sources. 2) Human-annotated datasets such as CharacterEval (Tu et al., 2024) and CharacterDial (Zhou et al., 2023) offer improved quality, but are expensive and difficult to scale. 3) Several efforts extract authentic dialogues from fictional works, such as ChatHaruhi (Li et al., 2023) and HPD (Chen et al., 2023). However, they rely on human efforts for individual sources and are hence hard to scale as well. 4) Furthermore, existing datasets offer limited representations and forms, i.e., mainly consisting of two-character or user-character question-answer pairs. These datasets support various purposes, including prompting, training, retrieval augmentation, and evaluation of RPLAs.
@@ 4.3 GCA Evaluation
Specifically, we employ LLM critics  [^4] This paper uses GPT-4o as the critic LLM by default. to identify flaw instances $\mathcal{F}$ in $\bar{M}$ of specific rubrics, such as “deviate from the original conversation” or “lack initiative and goals”, instead of directly outputting a score in previous work (Wang et al., 2024a; Tu et al., 2024). Each flaw $f$ is assigned a severity $v_{f}$ from 1 (minor) to 5 (severe). The initial score for each dimension is calculated as $s=100-\sum_{f\in\mathcal{F}}v_{f}$.
The rubrics are derived from human-annotated issues in extensive human-RPLA conversations from real users and our multi-agent simulations. We also consider evaluation dimensions from previous work (Shanahan et al., 2023; Shao et al., 2023; Chen et al., 2024a). For more informed evaluation, LLM critics are provided with additional materials, i.e., the original conversation $M$ and plot summary, besides data available to actor LLMs. Each dimension is assessed in independent LLM requests. The evaluation dimensions and their rubrics are summarized as follows:
@@
As longer simulations naturally make more flaws, we implement length correction to reduce bias in LLM judging following  Dubois et al. (2024). Specifically, we obtain the length corrected score as $s=100-\sum_{f\in\mathcal{F}}v_{f}+\lambda|\bar{M}|$, where $\lambda$ is set to 1.5 based on statistical analysis in §C.3. For detailed prompts and rubrics, please refer to §F.
@@ 5.1 Settings
Metrics  We report LLM-judged scores for each dimension, and their average as the overall score. For analysis, we also evaluate two traditional metrics based on N-gram, i.e., BLEU (Papineni et al., 2002) and ROUGE-L(Lin and Och, 2004) compared against original dialogues. Additionally, we report win rates versus GPT-3.5 and GPT-4o in §E.
Models  Our experiments cover numerous LLMs: 1) Close models, including Minimax Abab7-preview  [^5] https://www.minimaxi.com/, Doubao-pro  [^6] https://team.doubao.com/en/ Version:241215, Step-1-Flash and Step-2  [^7] https://www.stepfun.com/ Version: 241111 for Step-1-Flash, 241223 for Step-2 (internal)., GPT-3.5 (OpenAI, 2022), GPT-4o and GPT-4o Mini  [^8] https://openai.com/index/hello-gpt-4o/ Version:240806, Gemini-1.5-Pro  [^9] https://deepmind.google/technologies/gemini/pro/, Claude-3-Haiku and Claude-3.5-Sonnet  [^10] https://www.anthropic.com/ Version: 20240307 for Claude-3-Haiku, 20240620 for Claude-3.5-Sonnet; 2) Open models, including CoSER 8B and 70B, LLaMA-3.1-Instruct 8B and 70B (Dubey et al., 2024), Qwen-2-Instruct 7B and 72B (Yang et al., 2024), Vicuna-13B-1.5 (Zheng et al., 2023), Mixtral-8x7B (Jiang et al., 2024), DeepSeek-V3 (Liu et al., 2024) and Higgs-Llama-3-70B  [^11] https://boson.ai/higgs-v2/.
@@ 5.2 Main Results
Conversation Continuation Enables More Controlled Evaluation  Table  15 shows experiment results when multi-agent systems start from the first $k=3$ original messages. In this setting, model obtain higher scores compared to simulations from scratch ($k=0$), with reduced performance gaps between different models, especially for BLEU and ROUGE-L results. For example, the average score gap between Qwen-2 72B and 7B decreases from 11.5% ($k=0$) to 8.8% ($k=3$). This improvement occurs because the $k=3$ original messages guide the story direction and language style, particularly benefiting smaller models that typically struggle with complex role-playing instructions.
@@ 5.3 Ablation Studies
Inner Thoughts Benefit Role-Playing Training  We train CoSER model variants without inner thoughts, and evaluate them on various benchmarks. Results in Tables  3 and  4 show that models trained without inner thoughts consistently underperform regular CoSER models, exhibiting the value of inner thoughts for role-playing training.
@@ 5.4 CoSER Dataset for Retrieval Augmentation
We evaluate the value of our comprehensive data types for retrieval augmentation on CoSER Test. We explore three retrieval sources for a specific character: dialogues in related conversations (Conv.), as well as experiences (Expr.) and raw text in related plots. The retrieval system is based on FAISS (Douze et al., 2024) with BGE-M3 (Chen et al., 2024b) for text embeddings. As shown in Fig. 17, we observe: 1) Models consistently benefit from characters’ retrieved experiences and conversations, especially for CoSER 70B; 2) However, raw text retrieval barely enhances LLMs’ performance. Detailed experimental settings and results are provided in §E and Table  17.
@@ 5.5 Case studies
We conduct case studies to analyze LLMs’ performance in GCA simulation. Several cases are presented in Tables 7 to 12, from which we observe that: 1) CoSER models, trained on authentic dialogues, communicate more naturally, closely aligning with human speech patterns. 2) CoSER models better recall character-related knowledge, such as the iconic line “Grown enough to be wed, wed enough to be bedded” by Lysa Arryn in Table 7. This is consistent with their high BLEU and ROUGE-L scores. 3) CoSER models better portray sophisticated thinking process of humans. For example, in Tables 9 to 12, CoSER 70B faithfully reproduces Cersei Lannister’s suppressed anger as depicted in the original conversations, while other models, including GPT-4o and Claude-3.5-Sonnet, resort to a stereotypical portrayal of her arrogance and pride.
@@ A.1 Statistics and Analysis
As shown in Table 5, CoSER dataset is extensive and comprehensive, encompassing dialogue data from 771 books and 17,966 distinct characters. The dataset includes 30,069 unique plots and 29,798 conversations. On average, each conversation consists of approximately 13.2 utterances, with the entire dataset comprising a total of 392,900 utterances.
@@
Our book selection is derived from the Best Books Ever list on Goodreads, a curated collection of globally acclaimed literary works. These novels have garnered widespread recognition and appreciation from readers worldwide. Table  6 presents a comprehensive list of the top 100 books from our selection.
@@ B.1 Training Samples
Towards general role-playing capabilities across diverse scenarios and applications, we dynamically generate role-playing instructions (system prompts) using varied phrasings, formats, and data types, as shown in Tables  21 and  22. We consider instructions entirely in natural language, as well as those formatted with special symbols (such as ###, $=$), randomly sampling different formats and various expressions for the same semantics. We consider various configurations of the available data. Each sample may include (50%) or exclude (50%) the following elements : 1) Profiles of other characters in this conversation; 2) Summaries of the relevant plot; 3) Inner thoughts within the messages.

### v2

@@ Abstract
Abstract: Role-playing language agents (RPLAs) have emerged as promising applications of large language models (LLMs). However, simulating established characters presents a challenging task for RPLAs, due to the lack of authentic character datasets and nuanced evaluation methods using such data. In this paper, we present CoSER, a collection of a high-quality dataset, open models, and an evaluation protocol towards effective RPLAs of established characters. The CoSER dataset covers 17,966 characters from 771 renowned books. It provides authentic dialogues with real-world intricacies, as well as diverse data types such as conversation setups, character experiences and internal thoughts. Drawing from acting methodology, we introduce given-circumstance acting for training and evaluating role-playing LLMs, where LLMs sequentially portray multiple characters in book scenes. Using our dataset, we develop CoSER 8B and CoSER 70B, i.e., advanced open role-playing LLMs built on LLaMA-3.1 models. Extensive experiments demonstrate the value of the CoSER dataset for RPLA training, evaluation and retrieval. Moreover, CoSER 70B exhibits state-of-the-art performance surpassing or matching GPT-4o on our evaluation and three existing benchmarks, i.e., achieving 75.80% and 93.47% accuracy on the InCharacter and LifeChoice benchmarks respectively. Our code, dataset and models are available at: https://github.com/Neph0s/CoSER.
@@ 1 Introduction
In this paper, we introduce CoSER, a collection of authentic character data, along with open state-of-the-art models and evaluation protocol based on such data, for Coordinating LLM-Based Persona Simulation of Established Roles. The CoSER dataset is sourced from narratives and dialogues in 771 renowned books, processed via our LLM-based pipeline. CoSER differs from existing datasets in two fundamental ways: 1) CoSER extracts authentic, multi-character dialogues from acclaimed literary works, in contrast to LLM-synthesized question-answer pairs in previous work. Hence, our dataset maintains high source fidelity while exhibiting greater quality and complexity. 2) CoSER incorporates comprehensive types of data, as shown in Fig.  1: i) Besides character profiles and dialogues, CoSER encompasses plot summaries, character experiences, and conversation backgrounds, supporting various purposes including prompting, retrieval, model training and evaluation. ii) Conversations in CoSER capture characters’ actions and internal thoughts beyond surface-level speech, enabling RPLAs to simulate sophisticated cognitive and behavioral processes of humans, such as “[I’m nervous, but we have to do this] (Takes a deep breath) Alright, we …”. We provide a clear comparison between CoSER and existing datasets in Table  7.
@@
    Results of extensive experiments demonstrate the significant value of our dataset for the training, retrieval and evaluation of RPLAs. Notably, the CoSER models achieve state-of-the-art performance across four benchmarks and human evaluation for RPLAs.
@@ 2 Related Work
Using LLMs to simulate human personas represents a pioneering research direction. Early studies develop prototypes of RPLAs for fictional characters (Shao et al., 2023) and multi-agent systems that simulate human society (Park et al., 2023), while exploring the nature (Shanahan et al., 2023) and potential limitations (Liu et al., 2024b; Deshpande et al., 2023) of LLM role-play.  Chen et al. (2024a) present a comprehensive survey of relevant research.
Specifically, an RPLA leverages LLMs to create a simulated persona $\pi_{c}$ that simulates a real character $c$ based on its persona data $\mathcal{D}_{c}$. Effective RPLAs require both comprehensive, high-quality data $\mathcal{D}_{c}$ and advanced role-playing LLMs. Among various persona types, we focus on RPLAs for established characters, which should faithfully align with their characters’ complex backgrounds and nuanced personalities.
Datasets for RPLAs  Persona data $\mathcal{D}_{c}$ describe the real persona $c$ through various representations, including profiles (Yuan et al., 2024), dialogues (Wang et al., 2024a), experiences (Li et al., 2023) and multimodal information (Dai et al., 2024), e.t.c.. As shown in Table  7, existing datasets have several limitations. 1) Many are synthesized via LLMs’ responses to general instruction sets (Wang et al., 2024a; Chan et al., 2024) or character-specific questions (Shao et al., 2023), such as RoleBench (Wang et al., 2024a). However, LLM-synthesized data compromise authenticity and fidelity to original sources. 2) Human-annotated datasets such as CharacterEval (Tu et al., 2024) and CharacterDial (Zhou et al., 2023) offer improved quality, but are expensive and difficult to scale. 3) Several efforts extract authentic dialogues from fictional works, such as ChatHaruhi (Li et al., 2023) and HPD (Chen et al., 2023). However, they rely on human efforts for individual sources and are hence hard to scale as well. 4) Furthermore, existing datasets offer limited representations and forms, i.e., mainly consisting of two-character or user-character question-answer pairs. These datasets support various purposes, including prompting, training, retrieval augmentation, and evaluation of RPLAs.
@@ 4.3 GCA Evaluation
Specifically, we employ LLM critics  [^4] This paper uses GPT-4o as the critic LLM by default. to identify flaw instances $\mathcal{F}$ in $\bar{M}$ of specific rubrics, such as “deviate from the original conversation” or “lack initiative and goals”, instead of directly outputting a score in previous work (Wang et al., 2024a; Tu et al., 2024). Each flaw $f$ is assigned a severity $v_{f}$ from 1 (minor) to 5 (severe). The initial score for each dimension is calculated as $s=100-5*\sum_{f\in\mathcal{F}}v_{f}$.
Figure 3: We divide the RPLA evaluation into four quadrants (dimensions). The X-axis represents the evaluation perspective: Referenced (comparing with data in the source book) versus Inherent (assessing standalone quality). The Y-axis indicates the evaluation scope: Individual Agent versus Overall Simulation.
We consider four evaluation dimensions, as shown in Fig.  3. The rubrics are derived from human-annotated issues in extensive human-RPLA conversations from real users and our multi-agent simulations, as well as dimensions from previous work (Shanahan et al., 2023; Shao et al., 2023; Chen et al., 2024a). For more informed evaluation, LLM critics are provided with additional materials, i.e., the original conversation $M$ and plot summary, besides data available to actor LLMs. Each dimension is assessed in independent LLM requests. The evaluation dimensions and their rubrics are summarized as follows:
@@
As longer simulations naturally make more flaws, we implement length correction to reduce bias in LLM judging following  Dubois et al. (2024). Specifically, we obtain the length corrected score as $s=100-5*\sum_{f\in\mathcal{F}}v_{f}+\lambda|\bar{M}|$, where $\lambda$ is set to 1.5 based on statistical analysis in §C.3. For detailed prompts and rubrics, please refer to §G.
@@ 5.1 Settings
Metrics  We report LLM-judged scores for each dimension, and their average as the overall score. For analysis, we also evaluate two traditional metrics based on N-gram, i.e., BLEU (Papineni et al., 2002) and ROUGE-L(Lin and Och, 2004) compared against original dialogues. Additionally, we report win rates versus GPT-3.5 and GPT-4o in §D.
Models  Our experiments cover numerous LLMs: 1) Close models, including Minimax Abab7-preview  [^5] https://www.minimaxi.com/, Doubao-pro  [^6] https://team.doubao.com/en/ Version:241215, Step-1-Flash and Step-2  [^7] https://www.stepfun.com/ Version: 241111 for Step-1-Flash, 241223 for Step-2 (internal)., GPT-3.5 (OpenAI, 2022), GPT-4o and GPT-4o Mini  [^8] https://openai.com/index/hello-gpt-4o/ Version:240806, Gemini-1.5-Pro  [^9] https://deepmind.google/technologies/gemini/pro/, Claude-3-Haiku and Claude-3.5-Sonnet  [^10] https://www.anthropic.com/ Version: 20240307 for Claude-3-Haiku, 20240620 for Claude-3.5-Sonnet; 2) Open models, including CoSER 8B and 70B, LLaMA-3.1-Instruct 8B and 70B (Dubey et al., 2024), Qwen-2-Instruct 7B and 72B (Yang et al., 2024), Vicuna-13B-1.5 (Zheng et al., 2023), Mixtral-8x7B (Jiang et al., 2024), DeepSeek-V3 (Liu et al., 2024a) and Higgs-Llama-3-70B  [^11] https://boson.ai/higgs-v2/.
@@ 5.2 Main Results
Conversation Continuation Enables More Controlled Evaluation  Table  14 shows experiment results when multi-agent systems start from the first $k=3$ original messages. In this setting, model obtain higher scores compared to simulations from scratch ($k=0$), with reduced performance gaps between different models, especially for BLEU and ROUGE-L results. For example, the average score gap between Qwen-2 72B and 7B decreases from 11.5% ($k=0$) to 8.8% ($k=3$). This improvement occurs because the $k=3$ original messages guide the story direction and language style, particularly benefiting smaller models that typically struggle with complex role-playing instructions.
@@
5.3 Human Evaluation
We conduct human evaluation to further validate the performance of CoSER models and the reliability of LLM critics in GCA evaluation.
Annotations
We randomly select 60 samples from CoSER Test, and evaluate 7 representative LLMs (shown in Table 4) based on their simulation results. Three annotators independently score the LLMs on 20 samples each, using a 1 to 10 scoring scale, given the background context and original conversations. During this process, the annotators note that 1) GCA simulation deeply reflects LLMs’ role-playing abilities, compared with previous single-turn benchmarks; 2) Manual evaluation is highly challenging and time-consuming. It requires careful learning of complex background and abundant dialogues, and takes 15 minutes to evaluate 7 models for one sample on average. We report models’ average scores and win rate against other models.
Model Score Win Rate (%)
GPT-3.5 3.117 10.6
LLaMA-3.1-8B 3.600 19.4
Abab7-preview 4.533 37.5
CoSER-8B 4.567 38.6
GPT-4o 4.967 47.2
Claude-3.5-Sonnet 6.200 73.9
CoSER-70B 6.783 86.9
Table 4: Human evaluation on LLMs’ GCA simulation.
Results and Analysis
According to Table 4, we find that: 1) The results generally align with GCA evaluation, validating the reliability of our LLM-based evaluation protocol; 2) CoSER 70B maintain superior performance, obtaining the highest score (6.783) and win rate (86.9%); 3) Human evaluators show less preference for GPT models compared to GPT-4o as LLM judges, which likely stems from self-preference bias in LLM judges (Wataoka et al., 2024).
Alignment
We further study the alignment between LLM and human judges. We evaluate three judge models, GPT-4o, DeepSeek-V3 and DeepSeek-R1 (Guo et al., 2025). We consider standard GCA evaluation and ablation variants that remove: (i) original dialogues as reference, (ii) scoring rubrics, (iii) length correction, and (iv) dimension separation. We also compare with BLEU and ROUGE-L. The alignment is measured via model-wise preferences (§C.4).
The results in Table 5 reveal that: 1) All components of GCA evaluation contribute to improved alignment; 2) Reasoning models excel as judges, with DeepSeek-R1 achieving the highest alignment; 3) BLEU and ROUGE-L remain effective indicators for role-playing evaluation.
Scoring Method Alignment
GCA (GPT-4o) 68.6
    w/o reference 64.3
    w/o rubrics 65.1
    w/o length correction 64.5
    w/o dimension separation 65.2
GCA (DeepSeek-V3) 65.1
GCA (DeepSeek-R1) 77.5
    w/o reference 77.2
BLEU 75.3
ROUGE-L 72.0
Table 5: Alignment (%) between human judges and LLM judges (or automatic metrics).
@@ 5.4 Ablation Studies
Inner Thoughts Benefit Role-Playing Training  We train CoSER model variants without inner thoughts, and evaluate them on various benchmarks. Results in Tables  3 and  6 show that models trained without inner thoughts consistently underperform regular CoSER models, exhibiting the value of inner thoughts for role-playing training.
@@ 5.5 CoSER Dataset for Retrieval Augmentation
We evaluate the value of our comprehensive data types for retrieval augmentation on CoSER Test. We explore three retrieval sources for a specific character: dialogues in related conversations (Conv.), as well as experiences (Expr.) and raw text in related plots. The retrieval system is based on FAISS (Douze et al., 2024) with BGE-M3 (Chen et al., 2024b) for text embeddings. As shown in Fig. 16, we observe: 1) Models consistently benefit from characters’ retrieved experiences and conversations, especially for CoSER 70B; 2) However, raw text retrieval barely enhances LLMs’ performance. Detailed experimental settings and results are provided in §D and Table  16.
@@ 5.6 Case Studies
We conduct case studies to analyze LLMs’ performance in GCA simulation. Several cases are presented in Tables 17 to 22, from which we observe that: 1) CoSER models, trained on authentic dialogues, communicate more naturally, closely aligning with human speech patterns. 2) CoSER models better recall character-related knowledge, such as the iconic line “Grown enough to be wed, wed enough to be bedded” by Lysa Arryn in Table 17. This is consistent with their high BLEU and ROUGE-L scores. 3) CoSER models better portray sophisticated thinking process of humans. For example, in Tables 19 to 22, CoSER 70B faithfully reproduces Cersei Lannister’s suppressed anger as depicted in the original conversations, while other models, including GPT-4o and Claude-3.5-Sonnet, resort to a stereotypical portrayal of her arrogance and pride.
@@ A.1 Statistics and Analysis
As shown in Table 7, CoSER dataset is extensive and comprehensive, encompassing dialogue data from 771 books and 17,966 distinct characters. The dataset includes 30,069 unique plots and 29,798 conversations. On average, each conversation consists of approximately 13.2 utterances, with the entire dataset comprising a total of 392,298 utterances.
@@
Our book selection is derived from the Best Books Ever list on Goodreads, a curated collection of globally acclaimed literary works. These novels have garnered widespread recognition and appreciation from readers worldwide. Table  8 presents a comprehensive list of the top 100 books from our selection.
@@ A.4 Comparison with Existing Methods for Character Profiling
A.5 Safety
We conduct safety checks on the dataset and remove potentially problematic content. Specifically, we truncate 110 sensitive conversations and remove a total of 602 messages.
A.6 Comparison with PersonaHub
Recently, several studies have explored LLM role-playing using synthesized personas rather, such as PersonaHub (Chan et al., 2024). We provide a systematic comparison between CoSER and PersonaHub below.
Goal and Focus
The fundamental objectives of the two datasets differ significantly:
Aspect CoSER PersonaHub
Target Simulate established personas with high fidelity Synthesize instruction data for knowledge distillation
Persona Focus Depth and richness of character data Breadth of persona types
Key Capabilities Anthropomorphism, character fidelity, multi-character interaction General instruction following
Table 9: Comparison between CoSER and PersonaHub regarding goal and focus.
Dataset Quality and Design
The data sources and design reflect distinct research priorities:
Aspect CoSER PersonaHub
Data Source Authentic book dialogues GPT-4o synthesized conversations
Persona Origin Book characters LLM-synthesized personas
Persona Data Comprehensive profiles, dialogues, experiences Concise profiles with limited context
Design Priority Quality, authenticity, and character depth Quantity and diversity
Table 10: Comparison between CoSER and PersonaHub regarding dataset quality and design.
This comparison highlights that while PersonaHub excels in generating diverse synthetic personas for broad instruction-following training, CoSER focuses on preserving the authenticity and depth of established characters for high-fidelity role-playing applications.
@@ B.1 Training Samples
Towards general role-playing capabilities across diverse scenarios and applications, we dynamically generate role-playing instructions (system prompts) using varied phrasings, formats, and data types, as shown in Tables  27 and  28. We consider instructions entirely in natural language, as well as those formatted with special symbols (such as ###, $=$), randomly sampling different formats and various expressions for the same semantics. We consider various configurations of the available data. Each sample may include (50%) or exclude (50%) the following elements : 1) Profiles of other characters in this conversation; 2) Summaries of the relevant plot; 3) Inner thoughts within the messages.